solve the following system of equations
6x+7y=4
6x+9y=12

Answer by LinnW(1048)

6x+7y=4
6x+9y=12
**************
-1(6x+7y=4) = -6x-7y=-4
**************
-6x-7y=-4
6x+9y=12
2y = 8
y = 4
Substitute 4 for y in 6x+7y=4
6x+7(4)=4
6x + 28 = 4
add -28 to each side
6x = -24
divide each side by 6
x = -4